OpenStudy (anonymous):

Given f(x) = –5x + 1, what is f(–2)?

OpenStudy (texaschic101):

f(x) = -5x + 1 f(-2) = -5(-2) + 1 f(-2) = 10 + 1 f(-2) = 11

OpenStudy (anonymous):

thank you so much

OpenStudy (texaschic101):

anytime :)